OpenStudy (anonymous):

Jordan is a manager of a car dealership. He has 3 professional car washers to clean the entire lot of cars, Jennifer, Arianna, and Matthew. Jennifer can wash all the cars in 14 hours. Arianna can wash all the cars in 11 hours. Matthew is new to the car dealership, so no one knows how long it will take him. Jordan assigns all of them to wash the cars together. Explain to Jordan how this task can tell him how long it would take Matthew to complete the task if he worked by himself. Use complete sentences. anyone???

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

The new guy will take some time near to 11 to 14 hours. So probably more. This is because the experienced guys are taking such hours and since he's new, he might take more.

OpenStudy (anonymous):

^^uh not sure if you're just swinging a guess here, but a mathematical outlook would be something like this: The rates of Both Jennifer and Arianna are known, so you can model them like 1/11+1/14=25/154 is how much they can do together per hour. If you let all three complete a task, then you should get some time t, which it took to complete. Set up an equation: 25t/154+M*t=1 and plug in that t. Solve for M, which is Mark's rate.

OpenStudy (anonymous):

okay how would i solve that equation though im confused

OpenStudy (anonymous):

...You don't have the information to solve it. Once jordan gets the data we are looking for (t: the time it takes all three to complete it), THEN you can solve it.